A pump driven at 1440 rev/min having a displacement of 12.5 ml/rev and a volumetric efficiency of 87% is used to supply fluid to a circuit with two cylinders. If the cylinder dimensions are 63 mm bore x 35 mm rod x 250 mm stroke, and 80 mm bore x 55 mm rod x 150 mm stroke, find the minimum cycle time for both cylinders to extend and retract fully.


1
Expert's answer
2021-12-21T00:19:01-0500

minimum cycle time "=\frac{ (231\times RPM) }{(60\times Net_{area} )} =\frac{ (231\times 1440) }{(60\times \frac{\pi}{4}80^2\times 55 )} = 4.8 \approx 5"